5 circuits in an engineering system are connected in series. The probability that anyone of them works is 0.85. Assuming that they operate independently, what is the probability that the system works.
Answer #1

P(that system works) =P(all of 5 circuits should work as they are connected in series)

=0.85*0.85*0.85*0.85*0.85=0.443705